# Online Hypergraph Matching

### v2

@@ 4.4 Bound for the last phase $T$
In order to be able to get a lower bound on $\ell(t,i)$, we now relate it to a process which is simpler to analyze, defined as follows on ×$\mathbb{N}\times\mathbb{Z}/2$:
@@
where $B(t,\frac{1}{2})$ is the binomial distribution with parameters $t$ and $\frac{1}{2}$ (see Figure 5 for an illustration). An important property of this function that will help us prove Theorem 4.1 is the following upper bound:
$\sum_{y=0}^{\infty}\left(1-\psi\left(t,\frac{y}{2}\right)\right)\leq 1+\frac{1}{2}\sqrt{t}\qquad\forall t\geq 1.$ (8)
$\xi(t,i):=a\;\psi\left(t,q_{t}-i\right)+b-\varepsilon t,$
where $a:=2-2\ln\Big{(}(e+1)/2)\Big{)}\approx 0.76$ and $b:=2\ln\Big{(}(e+1)/2\Big{)}-1\approx 0.24$. Here, $a$ and $b$ are chosen such that whenever the algorithm exactly hits the threshold for every edge, we have $\xi(t,t/2)=\ln((e+1)/2)=\ell(t,t/2)$ for all even $t$ and $\lim_{t\to\infty}\xi(t,x)=1=\lim_{t\to\infty}\ell(t,x)$ for all $x$.
@@
Figure 5: Plot of $\psi(t,y)$ for two different values of $t$, the horizontal axis represents $y\in\mathbb{Z}/2$.
Let us first state the properties we need from the $\psi$ function.
Claim 4.1.
@@
    For all $t$, we have: $\psi(t,0)=\frac{1}{2}$.
  4. 4.
  5. 5.
    For all $t$, we have $\sum_{y=0}^{\infty}(1-\psi(t,\frac{1}{2}y))\leq 1+\frac{1}{2}\sqrt{t}$.
Proof.
The first two statements follow directly from the definition. The third statement follows from the symmetry of $B(t,\frac{1}{2})$ around $\frac{t}{2}$. For the fourth statement, let $X\sim B(t,\frac{1}{2})$ and $Y\sim B(1,\frac{1}{2})$ be independent. Then, we have:
[…]
$=1+2\sqrt{\operatorname{Var}\left[X\right]}=1+\frac{1}{2}\sqrt{t}.$
We need one additional lemma before being able to get a lower bound for the loads of the nodes. As a reminder, $r_{t}$ is the number of active nodes on each side of the bipartition at the end of phase $t$ and $q_{t}:=(r_{t}+1)/2$. For every active node $u=\sigma_{t}(i)$, where $i\leq r_{t}$, let us define $d_{t}(u)=|i-q_{t}|$. This quantity measures the distance, in index, between node $u$ and the active node at position $q_{t}$.
@@
$d_{t}(u)\leq d_{t-1}(u)+\frac{1}{2}\quad\text{if }i<q_{t}\qquad{and}\qquad d_{t}(u)\leq d_{t-1}(u)-\frac{1}{2}\quad\text{if }i>q_{t}.$
@@
We will prove this statement by induction on $t$. For the base case, consider $t=1$. There are two possibilities, either the edge $(1,1)$ does not make it to the threshold, i.e. $12f(\ell_{1}^{(1)})<$ 1, in which case $r_{t}=0$, $q_{t}=0$, and the statement is then trivially satisfied. If the edge $(1,1)$ makes it to the threshold, then $2f(\ell_{1}^{(1)})=2f(\ell(1,1))\geq 1$, which is equivalent to $\ell(1,1)\geq\ln((e+1)/2)$ by definition of $f(x)=e^{x}/(e+1)$. Observe that in this case $r_{t}=q_{t}=1$, leading to
@@
Suppose now by induction that the statement holds for $t-1$, let $r_{t}$ be the active nodes at the end of phase $t$, let $q_{t}:=(r_{t}+1)/2$ and consider an arbitrary $i\in\{1,\dots,\lceil q_{t}\rceil\}$. Let us first consider the case where $i=q_{t}=(r_{t}+1)/2$, which can only occur when $r_{t}$ is odd. Observe that this means the edge $(\sigma_{t}(i),\sigma_{t}(i))$ belongs to the matching $\mathcal{M}_{i}^{(t)}$ and exceeds the threshold, i.e. $\ell(t,i)\geq\ln((e+1)/2)$. Using the fact that $\psi(t,0)=1/2$ for all $t$, $i=q_{t}$ and the exact same arguments as above, we get
@@
Consider now the case where $i<q_{t}$. Let $e=(u,v)=(\sigma_{t}(i),\sigma_{t}(r_{t}+1-i))\in\mathcal{M}_{i}^{(t)}$ and observe that $e$ exceeds the threshold, i.e. $f(\ell^{(t)}_{u})+f(\ell^{(t)}_{v})=f(\ell(t,i))+f(\ell(t,r_{t}+1-i))\geq 1$. Let us pick indices $j,k$ such that $u=\sigma_{t-1}(j)$ and $v=\sigma_{t-1}(r_{t-1}+1-k)$. Observe that:
@@
Otherwise, we have $(\sigma_{t-1}(k),v)\in\mathcal{M}_{i}^{(t-1)}$ and by using the fact that the algorithm is $\varepsilon$-threshold respecting, we get $f(\ell({t-1,k}))+f(\ell_{v}^{(t-1)})\leq 1+\varepsilon$. By using the inductive hypothesis $\ell({t-1,k})\geq\xi(t-1,k)$, we get
$f(\ell_{v}^{(t-1)})\leq 1+\varepsilon-f(\xi(t-1,k)).$
Since edge $e$ exceeds the threshold at the end of phase $t$, we have $f(\ell^{(t-1)}_{u}+x_{e}^{(t)})+f(\ell^{(t-1)}_{v}+x_{e}^{(t)})\geq 1$, which leads to
$x_{e}^{(t)}$ $\geq-\ln\Big{(}f(\ell_{u}^{(t-1)})+f(\ell_{v}^{(t-1)})\Big{)}\geq-\ln\Big{(}1+\varepsilon-f(\xi(t-1,k))+f(\ell_{u}^{(t-1)})\Big{)}$
$\geq f(\xi(t-1,k))-f(\ell_{u}^{(t-1)})-\varepsilon=f(\xi(t-1,k))-f(\ell(t-1,j))-\varepsilon,$
where we have used the fact that $f(x)=e^{x}/(e+1)$ is an increasing function and $\ln(1+x)\geq x$. Finally, since we have $h^{\prime}(\ell_{u}^{(t-1)})=f(\ell_{u}^{(t-1)})\geq h\Big{(}\ln((e+1)/2)\Big{)}=1/2$:
$\ell(t,i)$ $=\ell_{u}^{(t-1)}+x_{e}^{(t)}\geq\ell_{u}^{(t-1)}+h^{\prime}(\ell_{u}^{(t-1)})\Big{(}\xi(t-1,k)-\ell(t-1,j)\Big{)}-\varepsilon$
$=\frac{1}{2}\ell(t-1,j)+\frac{1}{2}\xi(t-1,k)-\varepsilon\geq\frac{1}{2}\xi(t-1,j)+\frac{1}{2}\xi(t-1,k)-\varepsilon$
@@
[…]
$=\frac{a}{2}\Big{(}\psi\Big{(}t-1,d_{t}(u)+\frac{1}{2}\Big{)}+\psi\Big{(}t-1,d_{t}(u)-\frac{1}{2}\Big{)}\Big{)}+b-\varepsilon t$
$=a\;\psi\left(t,q_{t}-i\right)+b-\varepsilon t$
@@
where we have used the inductive hypothesis in the second inequality and the fourth property of the $\psi$ function in the second to last inequality. ∎
@@
Consider the end of phase $T-1$. Observe that the nodes $\sigma_{T-1}(k)\in U_{i}$ and $\sigma_{T-1}(k)\in V_{i}$ for $k\in\{1,\dots,\lceil q_{T-1}\rceil\}$ form a vertex cover of the final matching $\mathcal{M}_{i}^{(T)}$. Because of the degree constraints, this means that the value the algorithm can gain on the last phase $T$ is at most twice the following expression:
@@
$=a\sum_{k=1}^{\lceil q_{T-1}\rceil}\Big{(}1-\;\psi(T-1,q_{T-1}-k)\Big{)}+\varepsilon(T-1)\lceil q_{T-1}\rceil$
$\leq\varepsilon\>T^{2}+a\sum_{y=-1}^{\infty}\left(1-\>\psi\left(T-1,\frac{y}{2}\right)\right)\leq\varepsilon\>T^{2}+a\left(2+\frac{1}{2}\sqrt{T-1}\right)=O(\sqrt{T})+\varepsilon\;O(T^{2}).$
The first equality uses the relation $1-b=a$, the second inequality is due to ${q_{T-1}}=(r_{T-1}+1)/2\leq T/2$ and the change of index $y:=2(q_{T-1}-k)$, while the last inequality is by (8) and $\xi(t,-1)\leq 1$ for all $t\geq 1$. ∎
@@ 4.5 Connecting the matching $\mathcal{M}^{(t)}$ to the online nodes
The way now in which we apply this construction is by partitioning the matching $\mathcal{M}^{(t)}$ into submatchings based on the load (or fractional degree) of the vertices, and applying Lemma 4.3 on each submatching separately. More precisely, let us fix $\eta:=|\mathcal{M}^{(t)}|^{-1/3}$ and $N:=\left\lceil 2/\eta\right\rceil$. We partition the edges of the matching $\mathcal{M}^{(t)}$ into $N^{2}$ submatchings as follows
@@
4.6 Bound for the first $T-1$ phases
Recall from Section 4.1 that our instance consists of $T$ phases. For ease of analysis, we will split the total value gained by the algorithm into the value gained in each phase. For an algorithm $\mathcal{A}$, let $\mathcal{V}^{(t)}(\mathcal{A})$ denote the value obtained by $\mathcal{A}$ in phase $t$. The next lemma upper bounds $\mathcal{V}^{(t)}(\mathcal{A})$ for a threshold-respecting algorithm $\mathcal{A}$, in terms of the loads of the offline nodes at the end of phase $t-1$. Recall that $W^{(t)}$ is the set of online nodes which arrive during phase $t$.
@@ 4.6 Bound for the first $T-1$ phases
$\Delta(i,j)$ $:=\max\left\{x:f\left(\frac{i-1}{N}+x\right)+f\left(\frac{j-1}{N}+x\right)\leq 1\right\}$
$=\ln\left(\frac{e+1}{\exp\left(\frac{i-1}{N}\right)+\exp\left(\frac{j-1}{N}\right)}\right).$
@@
$\mathcal{V}^{(t)}(\mathcal{A})$ $\leq\sum_{i=1}^{m}\sum_{(u,v)\in\mathcal{M}^{(t)}_{i}}\left(1-f(\ell^{(t-1)}_{u})-f(\ell^{(t-1)}_{v})\right)+15\,|\mathcal{M}^{(t)}|^{2/3}.$ (9)
Fix a component $i\in[m]$, and let $\mathcal{E}_{i}:=\Big{\{}e\in\mathcal{M}^{(t-1)}_{i}\mid f(\ell_{u}^{(t-1)})+f(\ell_{v}^{(t-1)})\geq 1\Big{\}}$ be the subset of edges in the matching $\mathcal{M}^{(t-1)}_{i}$ which exceed the threshold at the end of phase $t-1$. By the construction of $\mathcal{M}^{(t)}_{i}$ in Section 4.3, we know that its node set consists of the nodes incident to $\mathcal{E}_{i}$, in addition to two new fresh nodes whose load is 0 at the end of phase $t-1$. This allows us to expand the inner sum in (9) as:
@@
Let $z$ be the vector obtained by repeating this procedure on every edge $e\in\mathcal{M}^{(t)}$. Then, $\mathbbm{1}^{\top}z\leq\varepsilon tm$ as $|\mathcal{M}^{(t)}|\leq tm$. Moreover, observe that the algorithm which assigns $x-z$ in phase $t$ is threshold-respecting on $W^{(t)}$. Thus, we can apply Lemma 4.5 to obtain the desired upper bound on $\mathbbm{1}^{\top}(x-z)$. ∎
@@ 4.7 Putting everything together
In this section, we complete the proof of Theorem 1.1. Since we assumed that the algorithm is $\varepsilon$-threshold respecting in the first $T-1$ phases, we can apply Corollary 4.1 to upper bound the value obtained in the first $T-1$ phases as
@@
We prove that for every $t\in[T]$, there exists a hypergraph matching of size $tm$ at the end of phase $t$. Let $C_{i}$ be a component with bipartition $U_{i}=[T]$ and $V_{i}=[T]$. It suffices to show that there exists a graph matching ~$\widetilde{\mathcal{M}}^{(t)}_{i}$ with vertex set $[t]$ on each side. This is because ~$\widetilde{\mathcal{M}}^{(t)}:=\cup_{i=1}^{m}\widetilde{\mathcal{M}}^{(t)}_{i}$ can be extended to a hypergraph matching in $\mathcal{H}$ by our construction (see Lemma 4.3). Let $E^{(t)}_{i}\subseteq\mathcal{M}^{(t)}_{i}$ be the edges whose endpoints are not active at the end of phase $t$. Then, a simple inductive argument on $t\geq 1$ shows that $\cup_{s=1}^{t-1}E^{(s)}_{i}\cup\mathcal{M}^{(t)}_{i}$ is a graph matching with vertex set $[t]$ on each side (see Figure 3 for an example). ∎
@@
Hence, letting $m\to\infty$, picking $T=o(\sqrt{m})$ such that $T\to\infty$ and setting $\varepsilon=o(1/T)$, we conclude that the competitive ratio is upper bounded by $(e-1)/(e+1)$, thus finishing the proof of Theorem 1.1.

### v3

@@ 4.4 Bound for the last phase $T$
In order to be able to get a lower bound on $\ell(t,i)$, we now relate it to a process which is simpler to analyze, defined as follows on $\mathbb{N}\times\mathbb{Z}/2$:
@@
where $B(t,\frac{1}{2})$ is the binomial distribution with parameters $t$ and $\frac{1}{2}$ (see Figure 6 for an illustration). This process has the following crucial properties, whose proofs are given in Appendix C.
@@
Figure 6: Plot of $\psi(t,y)$ for two different values of $t$, the horizontal axis represents $y\in\mathbb{Z}/2$. Here $\epsilon=0$, as it should be thought of as a very small constant.
Lemma 4.1.
@@
    For all $t$, we have $\sum_{y=0}^{\infty}(1-\psi(t,\frac{1}{2}y))\leq 1+\sqrt{t}$.
$\xi(t,i):=a\;\psi\left(t,q_{t}-i\right)+b-\varepsilon t.$
We choose $a$ and $b$ such that whenever the algorithm exactly hits the threshold for every edge, we have $\xi(t,t/2)+\epsilon t=\ln((e+1)/2)=\ell(t,t/2)$ for all even $t$ and $\lim_{t\to\infty}\xi(t,x)+\epsilon t=1=\lim_{t\to\infty}\ell(t,x)$ for all $x$. Hence, we set $a:=2-2\ln\Big((e+1)/2)\Big)\approx 0.76$ and $b:=2\ln\Big((e+1)/2\Big)-1\approx 0.24$. This choice will allow us to lower bound $\ell(t,i)$ by $\xi(t,i)$ for all $i\in\{1,\dots,\lceil q_{t}\rceil\}$ in Lemma 4.3.
We first need one more lemma. As a reminder, $r_{t}$ is the number of active nodes on each side of the bipartition at the end of phase $t$ and $q_{t}:=(r_{t}+1)/2$. For every active node $u=\sigma_{t}(i)$, where $i\leq r_{t}$, let us define $d_{t}(u)=|i-q_{t}|$, the distance of $u$ to the middle point $q_{t}$ of the active nodes at time $t$.
@@
[…]
@@
We will prove this statement by induction on $t$.
Base case. For the base case, consider $t=1$. There are two possibilities, either the edge $(1,1)$ does not make it to the threshold, i.e. $2f(\ell_{1}^{(1)})<1$, in which case $r_{t}=0$, $q_{t}=0$, and the statement is then trivially satisfied. If the edge $(1,1)$ makes it to the threshold, then $2f(\ell_{1}^{(1)})=2f(\ell(1,1))\geq 1$, which is equivalent to $\ell(1,1)\geq\ln((e+1)/2)$ by definition of $f(x)=e^{x}/(e+1)$. Observe that in this case $r_{t}=q_{t}=1$, leading to
@@
Inductive step. Suppose now by induction that the statement holds for t1t-1, let $r_{t}$ be the number of active nodes at the end of phase $t$, let $q_{t}:=(r_{t}+1)/2$ and consider an arbitrary $i\in\{1,\dots,\lceil q_{t}\rceil\}$. We consider the cases where $i=q_{t}$ and $i<q_{t}$ separately.
Let us first consider the case where $i=q_{t}=(r_{t}+1)/2$, which can only occur when $r_{t}$ is odd. Observe that this means the edge $(\sigma_{t}(i),\sigma_{t}(i))$ belongs to the matching $\mathcal{M}_{i}^{(t)}$ and exceeds the threshold, i.e. $\ell(t,i)\geq\ln((e+1)/2)$. Using the fact that $\psi(t,0)=1/2$ for all $t$, $i=q_{t}$ and the exact same arguments as above, we get
@@
Consider now the case where $i<q_{t}$. Let $e=(u,v)=(\sigma_{t}(i),\sigma_{t}(r_{t}+1-i))\in\mathcal{M}_{i}^{(t)}$ and observe that $e$ exceeds the threshold, i.e. $f(\ell^{(t)}_{u})+f(\ell^{(t)}_{v})=f(\ell(t,i))+f(\ell(t,r_{t}+1-i))\geq 1$. Let us pick indices $j,k$ such that $u=\sigma_{t-1}(j)$ and $v=\sigma_{t-1}(r_{t-1}+1-k)$. We have:
@@
Otherwise, we have $(\sigma_{t-1}(k),v)\in\mathcal{M}_{i}^{(t-1)}$ and by using the fact that the algorithm is $\varepsilon$-threshold respecting, we get $f(\ell({t-1,k}))+f(\ell_{v}^{(t-1)})\leq 1+\varepsilon$. By using the inductive hypothesis $\ell({t-1,k})\geq\xi(t-1,k)$, we get $f(\ell_{v}^{(t-1)})\leq 1+\varepsilon-f(\xi(t-1,k)).$ Since edge $e$ exceeds the threshold at the end of phase $t$, we have $f(\ell^{(t-1)}_{u}+x_{e}^{(t)})+f(\ell^{(t-1)}_{v}+x_{e}^{(t)})\geq 1$, which leads to
$x_{e}^{(t)}$ $\geq-\ln\Big(f(\ell_{u}^{(t-1)})+f(\ell_{v}^{(t-1)})\Big)$
$\geq-\ln\Big(1+\varepsilon-f(\xi(t-1,k))+f(\ell_{u}^{(t-1)})\Big)$ (by induction hypothesis)
$\geq f(\xi(t-1,k))-f(\ell_{u}^{(t-1)})-\varepsilon\qquad$ (since $\ln(1+x)\geq x$)
$\geq f^{\prime}(\ell_{u}^{(t-1)})\Big(\xi(t-1,k)-\ell_{u}^{(t-1)}\Big)-\varepsilon\qquad$ (by convexity of $f(x)=\frac{\exp(x)}{e+1}$)
$=f^{\prime}(\ell_{u}^{(t-1)})\Big(\xi(t-1,k)-\ell(t-1,j)\Big)-\varepsilon.$
Finally, since we have $f^{\prime}(\ell_{u}^{(t-1)})=f(\ell_{u}^{(t-1)})\geq f\Big(\ln((e+1)/2)\Big)=1/2$, so that $x_{e}^{(t)}\geq\frac{1}{2}\Big(\xi(t-1,k)-\ell(t-1,j)\Big)-\varepsilon$, and hence:
$\ell(t,i)$ $=\ell_{u}^{(t-1)}+x_{e}^{(t)}\geq\ell_{u}^{(t-1)}+\frac{1}{2}\Big(\xi(t-1,k)-\ell(t-1,j)\Big)-\varepsilon$
$=\frac{1}{2}\ell(t-1,j)+\frac{1}{2}\xi(t-1,k)-\varepsilon$
$\geq\frac{1}{2}\xi(t-1,j)+\frac{1}{2}\xi(t-1,k)-\varepsilon\quad$ (by induction hypothesis)
@@
$\geq\frac{a}{2}\Big(\psi\Big(t-1,d_{t}(u)+\frac{1}{2}\Big)+\psi\Big(t-1,d_{t}(v)-\frac{1}{2}\Big)\Big)+b-\varepsilon t\quad$ (by Lemma 4.2)
$=a\;\psi\left(t,q_{t}-i\right)+b-\varepsilon t\quad$ (by property (2) in Lemma 4.1)
@@
This concludes the proof of the inductive step and thus of the lemma. ∎
@@
Consider the end of phase T1T-1. Observe that the nodes $\sigma_{T-1}(k)\in U_{i}$ and $\sigma_{T-1}(k)\in V_{i}$ for $k\in\{1,\dots,\lceil q_{T-1}\rceil\}$ form a vertex cover of the final matching $\mathcal{M}_{i}^{(T)}$. Because of the degree constraints, this means that the value the algorithm can gain on the last phase $T$ is at most twice the following expression:
@@
$=a\sum_{k=1}^{\lceil q_{T-1}\rceil}\Big(1-\;\psi(T-1,q_{T-1}-k)\Big)+\varepsilon(T-1)\lceil q_{T-1}\rceil\quad$ (definition of $\xi$ and 1b=a1-b=a)
$\leq\varepsilon\>T^{2}+a\sum_{y=-1}^{\infty}\left(1-\>\psi\left(T-1,\frac{y}{2}\right)\right)\quad$ (by change of variables $y:=2(q_{T-1}-k)$)
$\leq\varepsilon\>T^{2}+a\left(2+\sqrt{T-1}\right)\qquad$ (by property (3) of Lemma 4.1)
$=O(\sqrt{T})+\varepsilon\;O(T^{2}).$
@@ 4.5 Connecting the matching $\mathcal{M}^{(t)}$ to the online nodes
The way now in which we apply this construction is by partitioning the matching $\mathcal{M}^{(t)}$ into submatchings based on the load (or fractional degree) of the vertices, and applying Lemma 4.4 on each submatching separately. More precisely, let us fix $\eta:=|\mathcal{M}^{(t)}|^{-1/3}$ and $N:=\left\lceil 2/\eta\right\rceil$.
We partition the edges of the matching $\mathcal{M}^{(t)}$ into $N^{2}$ submatchings as follows
@@
4.6 Bound for the first T1T-1 phases
Recall from Section 4.1 that our instance consists of $T$ phases. For ease of analysis, we will split the total value gained by the algorithm into the value gained in each phase. For an algorithm $\mathcal{A}$, let $\mathcal{V}^{(t)}(\mathcal{A})$ denote the value obtained by $\mathcal{A}$ in phase $t$. The next lemma upper bounds $\mathcal{V}^{(t)}(\mathcal{A})$ for a threshold-respecting algorithm $\mathcal{A}$, in terms of the loads of the offline nodes at the end of phase t1t-1. Recall that $W^{(t)}$ is the set of online nodes which arrive during phase $t$.
@@ 4.6 Bound for the first T1T-1 phases
$\Delta(i,j)$ $:=\max\left\{x:f\left(\frac{i-1}{N}+x\right)+f\left(\frac{j-1}{N}+x\right)\leq 1\right\}=\ln\left(\frac{e+1}{\exp\left(\frac{i-1}{N}\right)+\exp\left(\frac{j-1}{N}\right)}\right).$
@@
[…]
Fix a component $i\in[m]$, and let $\mathcal{E}_{i}:=\Big\{e\in\mathcal{M}^{(t-1)}_{i}\mid f(\ell_{u}^{(t-1)})+f(\ell_{v}^{(t-1)})\geq 1\Big\}$ be the subset of edges in the matching $\mathcal{M}^{(t-1)}_{i}$ which exceed the threshold at the end of phase t1t-1. By the construction of $\mathcal{M}^{(t)}_{i}$ in Section 4.3, we know that its node set consists of the nodes incident to $\mathcal{E}_{i}$, in addition to two new fresh nodes whose load is 0 at the end of phase t1t-1. This allows us to expand the inner sum in (8) as:
@@
Let $z$ be the vector obtained by repeating this procedure on every edge $e\in\mathcal{M}^{(t)}$. Then, $\mathbbm{1}^{\top}z\leq\varepsilon tm$ as $|\mathcal{M}^{(t)}|\leq tm$. Moreover, observe that the algorithm which assigns xzx-z in phase $t$ is threshold-respecting on $W^{(t)}$. Thus, we can apply Lemma 4.6 to obtain the desired upper bound on $\mathbbm{1}^{\top}(x-z)$.
@@ 4.7 Putting everything together
In this section, we complete the proof of Theorem 1.1. Since we assumed that the algorithm is $\varepsilon$-threshold respecting in the first T1T-1 phases, we can apply Corollary 4.1 to upper bound the value obtained in the first T1T-1 phases as
@@
We prove that for every $t\in[T]$, there exists a hypergraph matching of size $tm$ at the end of phase $t$. Let $C_{i}$ be a component with bipartition $U_{i}=[T]$ and $V_{i}=[T]$. It suffices to show that there exists a graph matching $\widetilde{\mathcal{M}}^{(t)}_{i}$ with vertex set $[t]$ on each side. This is because $\widetilde{\mathcal{M}}^{(t)}:=\cup_{i=1}^{m}\widetilde{\mathcal{M}}^{(t)}_{i}$ can be extended to a hypergraph matching in $\mathcal{H}$ by our construction (see Lemma 4.4). Let $E^{(t)}_{i}\subseteq\mathcal{M}^{(t)}_{i}$ be the edges whose endpoints are not active at the end of phase $t$. Then, a simple inductive argument on $t\geq 1$ shows that $\cup_{s=1}^{t-1}E^{(s)}_{i}\cup\mathcal{M}^{(t)}_{i}$ is a graph matching with vertex set $[t]$ on each side (see Figure 4 for an example). ∎
@@
Hence, letting $m\to\infty$, picking $T=o(\sqrt{m})$ such that $T\to\infty$ and setting $\varepsilon=o(1/T)$, we conclude that the competitive ratio is upper bounded by (e1)/(e+1)(e-1)/(e+1), thus finishing the proof of Theorem 1.1.